How would I simplify this expression 1/x+1 = 1/x-1? This is what I have x+1 for my answer.
Answer by Alan3354(69443)

How would I simplify this expression 1/x+1 = 1/x-1? This is what I have x+1 for my answer.
==============
It's an equation, not an expression. See the equal sign?
So you don't simplify it, you solve it.
1/x+1 = 1/x-1
Cross multiply.
x-1 = x+1
-1 = +1
no solution.